# BTGenBot: Behavior Tree Generation for Robotic Tasks with Lightweight LLMs

### v1

@@ IV-D Validation
You will be provided a summary of a task performed by a robot, and your objective is to express this task as a behavior tree in XML format. The behavior tree represents a mobile robot tasked to visit two locations: (7,1) and (4,8). The available actions are: "MoveTo" <root BTCPP_format="4">     <BehaviorTree>         <Sequence>             <MoveTo x="7" y="1"/>             <MoveTo x="4" y="8"/>         </Sequence>      </BehaviorTree> </root> The behavior tree represents a mobile robot tasked to visit a sequence of locations: ((0,0), (2,3), (4, 7), (5, 11)). The available actions are: "MoveTo"
@@
  <root BTCPP_format="4">   <BehaviorTree>   <Sequence> -   <Fallback>       <Sequence> -       <Action ID="moveToNewConfiguration" -               goal_pose="{goal_pose}" /> +       <Action ID="moveToNewConfiguration" />           <Sequence> -           <Action ID="CheckForTarget" -                   target="{target}" /> +           <Action ID="CheckForTarget" /> -             <Fallback>                 <Sequence> -                 <Action ID="ApproachTarget" -                         target="{target}" /> +                 <Action ID="ApproachTarget"/>                 </Sequence> -               <Action ID="MoveBack" /> -             </Fallback>           </Sequence>       </Sequence> -       <Action ID="MoveBack" /> -   </Fallback>   </Sequence>   </BehaviorTree>   </root>
@@ V Conclusion
We introduce a novel approach in robotics, enabling large language models with at most 7 billion parameters to generate executable behavior trees tailored for robot behaviors. We conducted a comparative analysis using Alpaca, LlamaChat, and CodeLlama, to identify the most suitable model for the task of generating ready-to-use behavior trees. To enhance the generation capabilities of these base models, we created a new instruction-following dataset specifically designed for fine-tuning behavior tree generation. The generated behavior trees have been evaluated in terms of syntactic accuracy, assessed with Groot2, and semantic accuracy, using our custom-developed validator. Furthermore, their performance was assessed within a simulation environment and in real-world deployment on a physical robot. Our assessment shows how fine-tuned models perform better than base models. In particular, LlamaChat is the best model overall, although CodeLlama still demonstrated notable performance. This work showcases the efficiency of compact large language models in directing autonomous agents like robots. Furthermore, behavior trees offer a flexible solution thanks to their inherent modularity, enabling seamless expansion through the incorporation of additional prompts and hence additional robot functionalities. Finally, our approach offers a plug-and-play solution, to transition from the LLM to direct execution on the robot.
The main limitation of using an LLM to generate robot behavior is the challenging tasks of evaluating and validating the correctness of the solution. In our work we managed to do so using a custom-developed validator, however, it is limited to behaviors where the solution is already known. In the future, we plan to deploy our system directly on the robot, therefore, we will focus on how to provide an automatic validation of the generated behavior tree. An option can be the use of milestones in the task to provide intermediate constraints and validate simpler behaviors. Alternatively, an auxiliary LLM can be used to regenerate the description from the XML and compare it with the original prompt.
The final aim of this work is to provide a direct interface between a user and the robot. Using this interface, the user can describe a task using natural language and the robot will execute it without the need for direct support. Solving this challenge has significant potential to improve human-robot interaction and the capabilities of autonomous robots in a variety of fields such as service robotics, autonomous industrial inspections, and last-mile logistics.

### v2

@@ IV-E Validation
{mdframed}
[linecolor=light-gray] You will be provided a summary of a task performed by a robot, and your objective is to express this task as a behavior tree in XML format.
The behavior tree represents a mobile robot tasked to visit two locations: (7,1) and (4,8). The available actions are: "MoveTo" {mdframed}[backgroundcolor=highlight-example, innertopmargin=-4pt, innerbottommargin=-4pt, innerleftmargin=2pt, innerrightmargin=2pt]
<root BTCPP_format="4">
<BehaviorTree>
<Sequence>
<MoveTo x="7" y="1"/>
<MoveTo x="4" y="8"/>
</Sequence>
</BehaviorTree>
</root>
The behavior tree represents a mobile robot tasked to visit a sequence of locations: ((0,0), (2,3), (4, 7), (5, 11)). The available actions are: "MoveTo"
@@
{mdframed}
[linecolor=light-gray]
<root BTCPP_format="4">
<BehaviorTree>
<Sequence>
- <Fallback>
<Sequence>
- <Action ID="moveToNewConfiguration"
- goal_pose="{goal_pose}" />
+ <Action ID="moveToNewConfiguration" />
<Sequence>
- <Action ID="CheckForTarget"
- target="{target}" />
+ <Action ID="CheckForTarget" />
- <Fallback>
<Sequence>
- <Action ID="ApproachTarget"
- target="{target}" />
+ <Action ID="ApproachTarget"/>
</Sequence>
- <Action ID="MoveBack" />
- </Fallback>
</Sequence>
</Sequence>
- <Action ID="MoveBack" />
- </Fallback>
</Sequence>
</BehaviorTree>
</root>
@@ V Conclusion
We introduce a novel approach in robotics, enabling large language models with at most 7 billion parameters to generate executable behavior trees for robots. We compared Alpaca, LlamaChat, and CodeLlama, to identify the most suitable model for generating ready-to-use behavior trees. To enhance the generation capabilities of these base models, we created a new instruction-following dataset specifically designed for fine-tuning behavior tree generation. The generated behavior trees have been evaluated in terms of syntactic accuracy, assessed with Groot2, and semantic accuracy, using a custom-developed validator. Furthermore, their performance was assessed in simulation and real-world deployment on a physical robot. Our assessment shows how fine-tuned models perform better than base models. In particular, LlamaChat is the best model overall, although CodeLlama still demonstrated notable performance. This work showcases the efficiency of compact large language models in directing autonomous agents like robots. Finally, our approach is a plug-and-play solution, to transition from the LLM to direct execution on the robot.
The main limitation of using an LLM to generate robot behaviors is the challenging task of evaluating and validating the correctness of the solution. In our work we managed to do so using a custom-developed validator, however, it is limited to behaviors with a known solution. In the future, to deploy our system directly on the robot, we will focus on the automatic validation of the generated behavior tree. An option can be the use of milestones in the task to provide intermediate constraints and validate simpler behaviors. Alternatively, an auxiliary LLM can be used to regenerate the description from the XML and compare it with the original prompt.
The final aim of this work is to provide a direct interface between a user and the robot. Using this interface, the user can describe a task using natural language and the robot will execute it autonomously. Solving this challenge has significant potential to improve human-robot interaction and the capabilities of robots in various fields such as service robotics, industrial inspections, and last-mile logistics.
Acknowledgment
This paper is supported by the FAIR (Future Artificial Intelligence Research) project, funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, investment 1.3, line on Artificial Intelligence).